The length of a rectangle is 3 times the width. If the width were increased by 3 meters while the length remained the same, the new rectangle would have an area of 30 sq meters. Find the length and width of the original rectangle.
Answer by mananth(16946)

The length of a rectangle is 3 times the width. If the width were increased by 3 meters while the length remained the same, the new rectangle would have an area of 30 sq meters. Find the length and width of the original rectangle.
...
let width be w
length = 3w
...
width w+3
length 3w
3w(w+3)=30
3w^2+9w-30=0
w^2+3w-10=0
(w+5)(w-2)=0
w=2 m the width
length = 3w= 6 m
...
m.ananth@hotmail.ca